Assume that your grandmother wants to give you generous gift. She wants you to choose which one of the following sets of cash flows you would like to receive:

Decision #1:   Which set of Cash Flows is worth more now?

 

Assume that your grandmother wants to give you generous gift.  She wants you to choose which one of the following sets of cash flows you would like to receive:

 

Option A:  Receive a one-time gift of $ 10,000 today.    

Option B:  Receive a $1400 gift each year for the next 10 years.  The first $1400 would be

     received 1 year from today.             

Option C:  Receive a one-time gift of $17,000 10 years from today.

 

Compute the Present Value of each of these options if you expect the interest rate to be 3% annually for the next 10 years.    Which of these options does financial theory suggest you should choose?

 

       Option A would be worth $__________ today.

       Option B would be worth $__________ today.

       Option C would be worth $__________ today.

       Financial theory supports choosing Option _______
Answer #1

Option with highest present value today should be chosen

Option A = 10,000 = $10,000

Option B = 1400*PVAF(3%, 10 years) = 1400*8.5302= $11,942.28

Option C = 17,000*PVF(3%, 10 years) = 17,000*0.7441 = $12,649.7

Option C should be chosen   


answered by: Miss. H